# Enhanced Single-Photon Detector: A framework for Superconducting-Level Performance without cryogenic cooling

### v2

@@ Abstract
High-performance single-photon detectors (SPDs) are indispensable components for a wide range of quantum optical applications. However, the reliance of state-of-the-art devices on superconducting materials imposes severe technological demands and necessitates challenging operational conditions, such as cryogenics, thereby hindering scalable implementation. To address this, we propose the Enhanced Single-Photon Detector (ESPD) framework, a novel paradigm for achieving high-performance SPDs through the iterative enhancement of conventional room-temperature SPDs. Relying entirely on non-superconducting components and eliminating cryogenic cooling requirements, the ESPD scheme can upgrade a legacy SPD, with detection efficiency (DE) about $59\%$ and dark count rate (DCR) of $10^{-2}$, to a device with superior performance metrics, achieving DE exceeding $93\%$ and DCR below $10^{-9}$. This performance rivals or even surpasses that of state-of-the-art superconducting SPDs, allowing the minimal tolerable channel transmission rate for quantum key distribution (QKD) protocols to be reduced by several orders of magnitude. Although the architecture requires substantial integration effort, the scheme can provide superconducting-level performance without cryogenic cooling, offering a clear path toward the widespread deployment of high-performance SPDs as well as related quantum technologies in infrastructure-constrained environments.
keywords:
Room-Temperature Single photon-detector, Detective efficiency, Dark count rate, Quantum optics, Quantum key distribution
1 Introduction
@@
Single-photon detectors (SPDs) play a pivotal role in quantum optics, as measurements are indispensable for information readout across virtually all applications. The two most critical performance metrics of an SPD are the detection efficiency (DE), denoted by $\eta$, and the dark count rate (DCR), denoted by $d$. The former affects the efficiency of nearly all quantum operations, while the latter imposes the ultimate limitation in high-attenuation scenarios. In terms of DE, numerous quantum experiments rely on SPD signals to retrieve results. In linear optical quantum computation, where success events are probabilistic and require the simultaneous detection of multiple photons, the success rate decreases exponentially with low-efficiency SPDs [1, 2, 3]. In Bell tests for quantum non-locality, verifying violations of the Clauser–Horne–Shimony–Holt or Eberhard inequalities routinely requires DE exceeding specific thresholds to close loopholes [4, 5, 6, 7, 8, 9, 10], such as approximately 67% [11] or 83% [12]. In heralded single-photon sources, where entangled pairs are generated, and one photon is detected to confirm the presence of the other, the efficiency of the SPD directly dictates the success rate [13, 14, 15, 16]. With respect to DCR, a primary constraint arises in quantum communication, particularly in quantum key distribution (QKD). The security of a practical QKD protocol relies on maintaining the quantum bit error rate (QBER) below a protocol-specific threshold, significantly lower than 50% [17, 18, 19, 20]. However, communication channels induce exponential photon attenuation [21, 22, 23, 24]. If the effective signal rate, determined by the channel transmission rate and thus communication distance, falls below the DCR of the SPDs, the QBER rises precipitously, as dark count events (which yield a 50% error rate) dominate the detection statistics. Consequently, the practical secure distance of QKD is strictly limited by the DCR of the SPDs [25, 26, 27].
@@
Driven by their critical importance, SPD has remained a continuous research focus since the inception of quantum technologies. While capable of operating at room temperature, traditional SPDs often suffer from both low DE and high DCR. For instance, a frequency upconversion SPD presented in 2004 operates at 300 K with DE about 59% and DCR of $10^{-2}$ [28], while an InGaAs/InP SPD presented in 2017 operates at 223 K with DE about 27.5% and DCR of $10^{-6}$ [29]. Such specifications are insufficient for advanced quantum tasks, which typically require the DE of SPD above 90% with DCR below $10^{-7}$. To suppress DCR and enhance DE, recent efforts have shifted toward designing SPDs using superconducting devices, such as superconducting nanowire SPDs (SNSPDs) [30, 31, 32, 33]. These SPDs achieve performance superior to traditional counterparts, boosting DE to over 90% while suppressing DCR by several orders of magnitude. This capability enables the implementation of demanding quantum tasks, such as long-distance QKD [19, 34, 35, 36]. A summary of recent representative advancements in superconducting SPDs is presented in Table 1.
@@
Table 1: Summary of recent representative performance benchmarks for state-of-the-art superconducting SPDs.
Year Researchers (et al.) Operation temperature Detection efficiency Dark count rate
@@
To overcome these limitations, we propose the enhanced SPD (ESPD) architecture that can meet advanced performance requirements, potentially outperforming superconducting ones, while relying solely on non-superconducting components suitable for room-temperature operation. The proposed method iteratively upgrades the performance of an existing conventional SPD level by level, ultimately to a high-performance one. As a demonstration of its capability, the ESPD scheme can enhance a non-superconducting SPD designed 20 years ago [28] with DE about 59% and DCR of $10^{-2}$ to one achieving DE exceeding 93% and DCR below $10^{-9}$, thereby can rival the state-of-the-art devices and reduce the minimal tolerable channel transmission rate in QKD by several orders of magnitude. Despite the substantial engineering integration effort required for its large-scale architecture, the ESPD method offers two key advancements: It not only provides a high-performance SPD design approach without cryogenic components, enabling a practical route toward accurate room-temperature photonic detection, but also establishes a paradigm shift for quantum information processing by leveraging multi-copy analysis and iterative enhancement to boost performance using conventional components. Consequently, this work opens a new avenue for the scalable deployment of high-performance SPDs and advances broader quantum optical technologies in infrastructure-constrained environments.
2 Result
@@
2.1 Design of the Enhanced Single-Photon Detector
@@
The design utilizes qubit logic. The schematic of the ESPD paradigm is illustrated in Fig. 1.
@@
2.2 Theoretical Formulations
2.2.1 Notations
@@
For convenience, assume that all controlled operations $C(F_{i})$ are identical, denoted simply as $C(F)$. Let $p$ be the transmission probability of the signal path through a single module (an (optional) state preparation followed by a $C(F)$). Specifically, $p$ is the probability that the output signal remains non-vacuum given a non-vacuum input of a single non-cascaded module, accounting for loss but irrespective of decoherence (since the state can be re-prepared). Let $P$ and $Q$ denote the probabilities of obtaining a non-vacuum output on the controlled path before detection, given that the input to the corresponding module was non-vacuum or vacuum, respectively. Hence, $Q$ represents the intrinsic error rate of the auxiliary path in the absence of $C(F)$ (i.e., errors from auxiliary state preparation and measurement (SPAM), but not including detection). We approximately assume $p$, $P$, and $Q$ are invariant across all $C(F)$ module operations.
@@
2.2.2 Assumptions
In practical scenarios, SPD DE typically ranges from 10% to 90%, while the DCR is generally below $10^{-3}$. Consequently, the hierarchy $0\leq d\ll\eta\leq 1$ holds. We further assume $d_{s}\ll P\eta_{s}$, which is trivial: If it were violated, the signal contribution from the controlled operations would become indistinguishable from device noise, rendering the operations practically insignificant. Additionally, we assume that the transmission rate of a single controlled module $p\approx 1$, SPAM error of the auxiliary path $Q\ll 1$, given the maturity of commercial optical components (see Section 3.1). Moreover, for analytical simplicity, we assume the controlled gates are implemented in a feedforward controlled configuration. This ensures that $n_{s+1}$ represents both the number of effective modules and the total physical modules required for level $s+1$. However, we emphasize that the generalized ESPD design remains valid for post-selection controlled gates (for a brief discussion, see Section 3.1.2). Finally, considering practical implementation costs, we assume $n_{s}$ and $k_{s}$ are moderate integers satisfying $1\leq k_{s}\leq n_{s}$.
2.2.3 Formulations
@@
$P_{s+1,k_{s+1},i}$ (4)
$=$ $(1-Q^{\prime}_{s+1})\sum_{j_{1}+j_{2}=k_{s+1}}^{n_{s+1}}\binom{i}{j_{1}}P_{s+1}^{j_{1}}(1-P_{s+1})^{i-j_{1}}\binom{n-i}{j_{2}}Q_{s+1}^{j_{2}}(1-Q_{s+1})^{n-i-j_{2}}$
@@
$P_{s+1,k_{s+1}}=$ $P^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$ (5)
$\qquad\qquad\qquad\qquad\qquad+(1-P^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$
$=$ $P^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}P_{s+1}^{k-1}(1-P_{s+1})^{n_{s+1}-k_{s+1}+1}$
$\qquad\qquad\qquad\qquad\qquad+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\quad+Q^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$
@@
2.3 Analytical Approximations and Stability Analysis
While Eq. (8) describes the exact dynamics, an analytical solution is generally intractable. However, employing the approximations in Eq.(1) to (3) and the assumptions from Section 2.2.2, we can derive simplified bounds to elucidate the system’s behavior.
@@ 2.3 Analytical Approximations and Stability Analysis
$\qquad\qquad\qquad\qquad\qquad+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$
@@
$=$ $x^{k-2}[\binom{n}{k-1}a(1-x)^{n-k}[(k-1)(1-x)-x(n-k+1)]$
$\qquad\qquad\qquad\quad+\sum_{j\geq k}\binom{n}{j}[jx^{j+1-k}(1-x)^{n-j}-x^{j-k+2}(1-x)^{n-j-1}(n-j)]]$
@@
$d_{s+1}\leq d_{s}\binom{n_{s+1}}{k_{s+1}-1}(Q+d_{s})^{k_{s+1}-1}(1-Q-d_{s})^{n_{s+1}-k_{s+1}+1}$ (12)
$\qquad\qquad\qquad\qquad\qquad\qquad\quad+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}(Q+d_{s})^{j}(1-Q-d_{s})^{n_{s+1}-j}$
@@
$d_{s+1}$ $\lesssim d_{s}\binom{n_{s+1}}{k_{s+1}-1}(Q+d_{s})^{k_{s+1}-1}+\binom{n_{s+1}}{k_{s+1}}(Q+d_{s})^{k_{s+1}}$ (13)
$=(Q+d_{s})^{k_{s+1}-1}[(d_{s}\binom{n_{s+1}}{k_{s+1}-1}+\binom{n_{s+1}}{k_{s+1}}(Q+d_{s})]$
@@
$\eta_{s+1}\geq$ $p^{n_{s+1}}[P^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$ (14)
$\qquad\qquad\qquad\qquad\qquad+(1-P^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}]$
$=$ $p^{n_{s+1}}[P^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}P_{s+1}^{k_{s+1}-1}(1-P_{s+1})^{n_{s+1}-k_{s+1}+1}$
$\qquad\qquad\qquad\qquad\qquad+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}]$
$\approx$ $p^{n_{s+1}}[\eta_{s}\binom{n_{s+1}}{k_{s+1}-1}(P\eta_{s})^{k_{s+1}-1}(1-P\eta_{s})^{n_{s+1}-k_{s+1}+1}$
$\qquad\qquad\qquad\qquad\qquad+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}(P\eta_{s})^{j}(1-P\eta_{s})^{n_{s+1}-j}]$
@@
2.4 Performance Simulations
While obtaining an exact analytical solution for the non-linear, time-variant dynamic system governing the ESPD in Eq.(8) is mathematically complex, numerical methods provide a robust and sufficient approach for practical design. In this section, we present numerical simulations to validate the ESPD paradigm, demonstrating that substantial performance gains are achievable even with moderate initial hardware specifications. For example, a three-level ESPD can transform a semiconductor SPD into one exhibiting DE $93\%$ and DCR $10^{-9}$, effectively removing the necessity for superconducting materials while matching or exceeding the performance of state-of-the-art SNSPDs.
Simulation Parameters: We evaluated the performance of the ESPD scheme using two distinct baseline SPDs. The first represents an older technology, characterized by low performance parameters $(\eta_{0},d_{0})=(59\%,10^{-2})$ [28], while the second represents more modern InGaAs/InP SPD performance, with $(\eta_{0},d_{0})=(27.5\%,10^{-6})$ [29]. Additionally, the controlled-gate parameters are set to $p=0.98$, which can be achieved by commercial devices [38], and $P=0.97$, achievable recently by feedforward C-NOT gate implementations in photonic platforms [39, 40, 41, 42], while, for the auxiliary path, the error rate is set to $Q=0.2\%$, justified by the extremely low SPAM errors (without detections) achievable across specific DOF and the inherent design freedom in selecting the appropriate DOF for the ESPD implementation (see Section 3.1). We also implemented additional simulations using degraded parameter settings to evaluate robustness (also see Section 3.1).
Performance Outcomes: Achievable performance metrics across varying configurations and baseline architectures are summarized in Table 2 and 3 with visual representations in Fig. 2 and 3. The results indicate that the ESPD scheme can significantly enhance the capabilities of conventional SPDs. Notably, with only three or four levels of enhancement, the detectors achieve performance metrics (DE $93\%$, DCR $10^{-9}$) that rival or surpass those of advanced superconducting SPDs.
(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
Table 2: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$.

### v3

# A Theoretical Framework for Iteratively Enhanced Room-Temperature Single-Photon Detection

@@ Abstract
High-performance photon detection is indispensable in a wide range of quantum-optical applications and is conventionally treated as a fixed device-level operation based on single-photon detectors (SPDs). However, state-of-the-art SPDs rely on superconducting materials, which impose severe technological demands and require challenging operational conditions such as cryogenic cooling, thereby hindering scalable implementation. Here, we propose the enhanced single-photon detector (ESPD) framework, a theoretical paradigm supported by numerical simulations, that reformulates single-photon detection as an iteratively enhanced process based on state preparation, controlled operations, projective measurements, and multi-copy analysis, and enables substantial performance improvement using exclusively room-temperature components. Numerical simulations indicate that, within a physically motivated parameter regime, the ESPD framework can upgrade a legacy SPD with a detection efficiency (DE) of about 59% and a dark count rate (DCR) of about $10^{-2}$ to effective performance metrics with DE exceeding 93% and DCR below $10^{-9}$, comparable to those of superconducting SPDs. As a consequence, the minimal tolerable channel transmission rate for quantum key distribution protocols can be reduced by several orders of magnitude. While its physical realization would require substantial experimental integration, the ESPD framework establishes a general system-level perspective on photon detection, highlighting the potential of iterative quantum processing for overcoming intrinsic detector limitations at room temperature.
preprint: APS/123-QED
I Introduction
@@
Single-photon detection plays a central role in quantum optics and underlies a wide range of quantum-optical applications. Conventionally, the detection process is treated as a fixed device-level operation implemented by single-photon detectors (SPDs). The two most critical performance metrics of an SPD are the detection efficiency (DE), denoted by $\eta$, and the dark count rate (DCR), denoted by $d$. The former affects the efficiency of nearly all quantum operations, while the latter imposes the ultimate limitation in high-attenuation scenarios. In terms of DE, numerous quantum experiments rely on SPD signals to retrieve results. In linear optical quantum computation, where success events are probabilistic and require the simultaneous detection of multiple photons, the success rate decreases exponentially with low-efficiency SPDs [1, 2, 3]. In Bell tests for quantum non-locality, verifying violations of the Clauser–Horne–Shimony–Holt or Eberhard inequalities routinely requires DE exceeding specific thresholds to close loopholes [4, 5, 6, 7, 8, 9, 10], such as approximately 67% [11] or 83% [12]. In heralded single-photon sources, where entangled pairs are generated, and one photon is detected to confirm the presence of the other, the efficiency of the SPD directly dictates the success rate [13, 14, 15, 16]. With respect to DCR, a primary constraint arises in quantum communication, particularly in quantum key distribution (QKD). The security of a practical QKD protocol relies on maintaining the quantum bit error rate (QBER) below a protocol-specific threshold, significantly lower than 50% [17, 18, 19, 20]. However, communication channels induce exponential photon attenuation [21, 22, 23, 24]. If the effective signal rate, determined by the channel transmission rate and thus communication distance, falls below the DCR of the SPDs, the QBER rises precipitously, as dark count events (which yield a 50% error rate) dominate the detection statistics. Consequently, the practical secure distance of QKD is strictly limited by the DCR of the SPDs [25, 26, 27].
@@
Driven by their critical importance, SPD has remained a continuous research focus since the inception of quantum technologies. While capable of operating at room temperature, traditional SPDs often suffer from both low DE and high DCR. For instance, a frequency upconversion SPD presented in 2004 operates at 300 K with DE about 59% and DCR of $10^{-2}$ [28], while an InGaAs/InP SPD presented in 2017 operates at 223 K with DE about 27.5% and DCR of $10^{-6}$ [29]. Such specifications are insufficient for advanced quantum tasks, which typically require the DE of SPD above 90% with DCR below $10^{-7}$.
To suppress DCR and enhance DE, recent efforts have shifted toward designing SPDs using superconducting devices, such as superconducting nanowire SPDs (SNSPDs) [30, 31, 32, 33]. These SPDs achieve performance superior to traditional counterparts, boosting DE to over 90% while suppressing DCR by several orders of magnitude. This capability enables the implementation of demanding quantum tasks, such as long-distance QKD [19, 34, 35, 36]. A summary of recent representative advancements in superconducting SPDs is presented in Table 1.
@@
Table 1: Summary of recent representative performance benchmarks for state-of-the-art superconducting SPDs. OT: Operation Temperature, DE: Detective efficiency, DCR: Dark count rate.
Year Researchers (et al.) OT DE DCR
@@
To address the limitations imposed by the suboptimal performance of room-temperature SPDs and cryogenic operation in superconducting alternatives, we propose the enhanced single-photon detector (ESPD) framework, a system-level theoretical paradigm supported by numerical simulations for improving single-photon detection using exclusively room-temperature components. The ESPD framework reformulates single-photon detection as an iteratively enhanced process rather than a fixed device-level operation, in which detection performance is progressively improved through the integration of state preparation, controlled operations, projective measurements, and multi-copy analysis. As a representative numerical benchmark, we show that, within a physically motivated parameter regime, the ESPD framework can enhance a legacy non-superconducting SPD reported two decades ago [28], upgrading an initial detector with DE of about 59% and DCR of $10^{-2}$ to effective performance metrics with DE exceeding 93% and DCR below $10^{-9}$. These values are comparable to those of state-of-the-art superconducting SPDs and relax the minimal tolerable channel transmission rate for quantum key distribution by several orders of magnitude. While the physical realization of such an architecture would require substantial experimental integration, the ESPD framework provides a general platform for the theoretical and numerical investigation of enhancement mechanisms in room-temperature photon detection, highlighting the way in which system-level quantum processing can overcome intrinsic device limitations.
II Method and Result
@@
II.1 Paradigm of the Enhanced Single-Photon Detection
@@
With qubit logic, the schematic of the ESPD paradigm is illustrated in Fig. 1.
@@
Physical intuition: Intuitively, each level of the ESPD framework increases the number of effective detection opportunities for an input signal, thereby enhancing the overall detection capability. In an idealized limit where all components except the SPDs are assumed to be perfect, the $(s+1)$-th level generates $n_{s+1}$ identical copies of the incoming signal by the controlled modules, which are independently detected by $n_{s+1}$ SPDs. A non-vacuum outcome is reported if at least $k_{s+1}$ positive detection events are registered. Under this idealized assumption, the detection efficiency can be enhanced by choosing a small threshold $k_{s+1}$, for example, when $k_{s+1}=1$, the effective DE becomes $1-(1-\eta_{s})^{n_{s+1}}$. Furthermore, the DCR is strongly suppressed, since a false positive outcome occurs only when dark-count events arise in at least $k_{s+1}$ detectors, leading to an approximate scaling $d_{s}^{k_{s+1}}$, which decreases dramatically when $k_{s+1}$ is increased. This physical intuition remains valid beyond the idealized limit. When imperfections in the auxiliary components introduce errors that are sufficiently small relative to the activation probability of a genuine input signal, the ESPD framework continues to improve detection performance. A rigorous formulation of this behavior, accounting for non-ideal devices and error propagation, is developed in Section II.2.
II.2 Theoretical Formulations
The rigorous theoretical formulation of the ESPD framework is developed in this subsection. We first introduce the notations and assumptions, and then present the exact analytical expressions.
II.2.1 Notations and Basic Assumptions
@@
For convenience, assume that all controlled operations $C(F_{i})$ are identical, denoted simply as $C(F)$, and they are implemented in a feedforward controlled configuration, which ensures that $n_{s+1}$ represents both the number of effective modules and the total physical modules required for level $s+1$. However, we emphasize that the generalized ESPD paradigm remains valid for post-selection controlled gates (for a brief discussion, see Section III.1.2). Let $p$ be the transmission probability of the signal path through a single module (an (optional) state preparation followed by a $C(F)$). Specifically, $p$ is the probability that the output signal remains non-vacuum given a non-vacuum input of a single non-cascaded module, accounting for loss but irrespective of decoherence (since the state can be re-prepared). Let $P$ and $Q$ denote the probabilities of obtaining a non-vacuum output on the controlled path before detection, given that the input to the corresponding module was non-vacuum or vacuum, respectively. Hence, $Q$ represents the intrinsic error rate of the auxiliary path in the absence of $C(F)$ (i.e., errors from auxiliary state preparation and measurement (SPAM), but not including detection). We approximately assume $p$, $P$, and $Q$ are invariant across all $C(F)$ module operations.
@@
II.2.2 Parameter Approximation Assumptions
The assumptions below are employed solely to derive analytical approximations in Section II.3, as well as the approximation part in Eq. (1) and (2). They are not required for the general theoretical formulation of the ESPD framework nor for the numerical simulations presented in this work. Although these assumptions entail relatively stringent requirements, they are physically motivated by realistic components and will be discussed in detail in Section III.1.
In practical scenarios, SPD DE typically ranges from 10% to 90%, while the DCR is generally below $10^{-3}$. Consequently, the hierarchy $0\leq d\ll\eta\leq 1$ holds. We further assume $d_{s}\ll P\eta_{s}$, which is trivial: If it were violated, the signal contribution from the controlled operations would become indistinguishable from device noise, rendering the operations practically insignificant. Additionally, we assume that the transmission rate of a single controlled module $p\approx 1$, SPAM error of the auxiliary path $Q\ll 1$, given the maturity of commercial optical components (see Section III.1). Finally, considering practical implementation costs, we assume $n_{s}$ and $k_{s}$ are moderate integers satisfying $1\leq k_{s}\leq n_{s}$.
II.2.3 Formulations
@@
$P_{s+1,k_{s+1},i}=$ $(1-Q^{\prime}_{s+1})\sum_{j_{1}+j_{2}=k_{s+1}}^{n_{s+1}}\binom{i}{j_{1}}P_{s+1}^{j_{1}}(1-P_{s+1})^{i-j_{1}}\binom{n-i}{j_{2}}Q_{s+1}^{j_{2}}(1-Q_{s+1})^{n-i-j_{2}}$ (4)
@@
$P_{s+1,k_{s+1}}=$ $P^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}+(1-P^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$ (5)
$=$ $P^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}P_{s+1}^{k-1}(1-P_{s+1})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$
@@
$+Q^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$
@@
II.3 Analytical Approximations and Stability Analysis
While Eq. (8) describes the exact dynamics, an analytical solution is generally intractable. However, employing the approximations in Eq.(1) to (3) and the assumptions from Section II.2.2, we can derive simplified bounds to elucidate the system’s behavior.
@@ II.3 Analytical Approximations and Stability Analysis
$+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$
@@
$=$ $x^{k-2}[\binom{n}{k-1}a(1-x)^{n-k}[(k-1)(1-x)-x(n-k+1)]+\sum_{j\geq k}\binom{n}{j}[jx^{j+1-k}(1-x)^{n-j}-x^{j-k+2}(1-x)^{n-j-1}(n-j)]]$
@@
$d_{s+1}\leq$ $d_{s}\binom{n_{s+1}}{k_{s+1}-1}(Q+d_{s})^{k_{s+1}-1}(1-Q-d_{s})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}(Q+d_{s})^{j}(1-Q-d_{s})^{n_{s+1}-j}$ (12)
@@
$d_{s+1}\lesssim d_{s}\binom{n_{s+1}}{k_{s+1}-1}(Q+d_{s})^{k_{s+1}-1}+\binom{n_{s+1}}{k_{s+1}}(Q+d_{s})^{k_{s+1}}=(Q+d_{s})^{k_{s+1}-1}[(d_{s}\binom{n_{s+1}}{k_{s+1}-1}+\binom{n_{s+1}}{k_{s+1}}(Q+d_{s})]$ (13)
@@
$\eta_{s+1}\geq$ $p^{n_{s+1}}[P^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}+(1-P^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}]$ (14)
$=$ $p^{n_{s+1}}[P^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}P_{s+1}^{k_{s+1}-1}(1-P_{s+1})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}]$
$\approx$ $p^{n_{s+1}}[\eta_{s}\binom{n_{s+1}}{k_{s+1}-1}(P\eta_{s})^{k_{s+1}-1}(1-P\eta_{s})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}(P\eta_{s})^{j}(1-P\eta_{s})^{n_{s+1}-j}]$
@@
II.4 Performance Simulations
Table 2: Performance of an ESPD initialized with a SPD from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$.